# Mitigating Hallucinations in Large Vision-Language Models via Summary-Guided Decoding

### v1

@@ 6 Related works
Mitigating Language Priors in LVLMs. Large Vision-Language models (LVLMs) extend pre-trained Large Language Models (LLMs) by incorporating visual tokens, enabling them to process visual content Liu et al. (2023); Dai et al. (2023); Zhu et al. (2023). In LVLM architectures, the language model is significantly larger than the vision model, creating an imbalanced structure where the language model exerts more significant influence. As a result of this imbalance, the model tends to rely on linguistic patterns rather than adequately considering the visual information provided, a phenomenon known as the language prior problem Guan et al. (2024); Lee et al. (2024a, b). To address this issue, several studies have explored contrastive decoding techniques to mitigate the model’s over-reliance on language priors. Visual Contrastive Decoding (VCD) Leng et al. (2023) works by utilizing distorted images, which amplify the language prior, and Instruction Contrastive Decoding (ICD) Wang et al. (2024) introduces misleading instructions to achieve a similar effect. Both methods aim to reduce the language prior’s dominance by leveraging these amplified conditions to adjust the model’s behavior. Additionally, Multi-Modal Mutual Information Decoding (M3ID) Favero et al. (2024) identified that as the token length increases, the model dilutes visual information, leading to a more substantial reliance on language priors. To counter this, M3ID applies more assertive contrastive decoding techniques as the token length grows to calibrate the model’s over-reliance on language priors. However, contrastive decoding can disrupt the distribution of tokens essential for language modeling, leading to a decline in text quality. Additionally, due to the language prior, the two output distributions being contrasted may become more similar, making it less effective in reducing hallucinations. Our method, Summary-Guided Decoding (SGD), addresses these issues by using summarization techniques to naturally reduces the influence of language priors, allowing the model to focus more on the image. Furthermore, SGD preserves text quality by controlling only the POS tokens relevant to the image.
@@ 7 Conclusion
In this paper, we introduce Summary-Guided Decoding (SGD) as a novel method to mitigate object hallucinations in LVLMs. Our analysis reveals that as token sequences grow, LVLMs tend to increasingly rely on language priors, reducing the influence of visual information during the decoding process. To address this, SGD employs summarization techniques to shorten token length, encouraging the model to incorporate more visual details while controlling only the image-related POS tokens to maintain text quality. Our experimental results demonstrate that SGD significantly reduces object hallucination and achieves an optimal balance between factual accuracy and recall in both short and long description tasks.
@@
In this paper, we propose a Summary-Guided Decoding (SGD) to mitigate object hallucinations in Large Vision-Language Models (LVLMs). However, this approach comes with some limitations.
@@
Second, while summarization effectively reduces input length and helps mitigate hallucinations, it may also result in the loss of critical contextual information. Additionally, excessively long summaries can increase LVLMs’ dependence on language priors, which may degrade the performance of SGD. Therefore, it is crucial for future work to train LVLMs in a way that inherently avoids over-reliance on language priors, even when token lengths are extended.
Lastly, we employ part-of-speech (POS) tagging to distinguish between image-related and language-related tokens. However, relying solely on POS tagging for this differentiation can be problematic. The development of more advanced methods for token distinction could enhance the effectiveness of SGD and create further synergies with this approach.
@@
In this paper, we contribute to the future development of a safe and reliable AI community by conducting research focused on reducing hallucinations in Large Vision-Language Models.
@@
However, the models we used may still exhibit biases inherent in the underlying datasets and training processes. While our focus was on biases related to language priors, we acknowledge the need to address other potential biases as well.
@@ Appendix A Details of CODE, Hyperparameters, and GPU Cost
We conducted our experiments based on the OPERA Huang et al. (2024) code base which is publicly available. We used the publicly available code provided by the authors for the VCD and OPERA methods. While M3ID and ICD were implemented from scratch due to the lack of public code. For VCD, OPERA, and ICD, we used the hyperparameters as specified in their respective papers. Since only LLAVA 1.5’s hyperparameters were reported in M3ID, we applied these hyperparameters to both LLAVA 1.5 and InstructBLIP for our experiments. Also, we set repetition penalty as 1. All the experiments are conducted using 1 NVIDIA RTX A6000 GPU.
@@ Appendix B Experimental Settings for Analyzing Language Priors
We generated descriptions using LLAVA 1.5 7B for 5000 images from the MSCOCO 2014 validation dataset Lin et al. (2015) and annotated each token to determine whether it represented an object hallucination, defined as an object not present in the image, using the CHAIR metric pipeline Rohrbach et al. (2019) for evaluation.
Appendix C Distilled Flan-T5-base model
First, we employed LLAVA 1.5 7B to perform Summary-Guided Decoding with Self-Summarization when generating descriptions for 5,000 images from the MSCOCO dataset. During this process, LLAVA 1.5 iteratively summarized the previous sentence, and we saved each previous sentence along with its corresponding summarized sentence as a pair. This paired dataset was then used to fine-tune the Flan-T5-base model with the prompt “What is a summary of this text?” for training purposes.
Appendix D Summarize Prompt for Summary-Guided decoding
In SGD, we used summary prompt as
‘‘USER: Summarize the following
@@ Appendix D Summarize Prompt for Summary-Guided decoding
Appendix E GPT-4o Prompt for text quality evaluation
@@
Appendix F CHAIR metric on various token length
@@
Appendix G Case Study
This is the case study of Summary-Guided Decoding (SGD), Visual Contrastive Decoding (VCD), and Multi-Modal Mutual Information Decoding (M3ID) in generating up to 256 tokens in detailed captioning task. Case study shows that SGD generated image-related words, while VCD showed a tendency to hallucinate by relying on the word ‘tie’ during decoding. Additionally, M3ID exhibited issues in language modeling.
Figure 9: SGD case study.

### v3

@@ 6 Related works
Mitigating Language Priors in LVLMs. Large Vision-Language models (LVLMs) extend pre-trained Large Language Models (LLMs) by incorporating visual tokens, enabling them to process visual content Liu et al. (2023); Dai et al. (2023); Zhu et al. (2023). In LVLM architectures, the language model is significantly larger than the vision model, creating an imbalanced structure where the language model exerts more significant influence. As a result of this imbalance, the model tends to rely on linguistic patterns rather than adequately considering the visual information provided, a phenomenon known as the language prior problem Guan et al. (2024); Lee et al. (2024b, c). Several studies have explored ways to control LLM outputs to better align with desired objectives Li et al. (2023b); Hallinan et al. (2023); Kim et al. (2023, 2024b). Similarly, research on LVLMs has focused on contrastive decoding techniques to reduce the model’s over-reliance on language priors Manevich and Tsarfaty (2024). Visual Contrastive Decoding (VCD) Leng et al. (2023) works by utilizing distorted images, which amplify the language prior, and Instruction Contrastive Decoding (ICD) Wang et al. (2024) introduces misleading instructions to achieve a similar effect. Both methods aim to reduce the language prior’s dominance by leveraging these amplified conditions to adjust the model’s behavior. Additionally, Multi-Modal Mutual Information Decoding (M3ID) Favero et al. (2024) identified that as the token length increases, the model dilutes visual information, leading to a more substantial reliance on language priors. To counter this, M3ID applies more assertive contrastive decoding techniques as the token length grows to calibrate the model’s over-reliance on language priors. However, contrastive decoding can disrupt the distribution of tokens essential for language modeling, leading to a decline in text quality. Additionally, due to the language prior, the two output distributions being contrasted may become more similar, making it less effective in reducing hallucinations. Our method, Summary-Guided Decoding (SumGD), addresses these issues by using summarization techniques to naturally reduces the influence of language priors, allowing the model to focus more on the image. Furthermore, SumGD preserves text quality by controlling only the POS tokens relevant to the image.
@@ 7 Conclusion
In this paper, we introduce Summary-Guided Decoding (SumGD) as a novel method to mitigate object hallucinations in LVLMs. Our analysis reveals that as token sequences grow, LVLMs tend to increasingly rely on language priors, reducing the influence of visual information during the decoding process. To address this, SumGD employs summarization techniques to shorten token length, encouraging the model to incorporate more visual details while controlling only the image-related POS tokens to maintain text quality. Our experimental results demonstrate that SumGD significantly reduces object hallucination and achieves an optimal balance between factual accuracy and recall in both short and long description tasks.
@@
In this paper, we propose a Summary-Guided Decoding (SumGD) to mitigate object hallucinations in Large Vision-Language Models (LVLMs). However, this approach comes with some limitations.
@@
Second, while summarization effectively reduces input length and helps mitigate hallucinations, it may also result in the loss of critical contextual information. Additionally, excessively long summaries can increase LVLMs’ dependence on language priors, which may degrade the performance of SumGD. Therefore, it is crucial for future work to train LVLMs in a way that inherently avoids over-reliance on language priors, even when token lengths are extended.
Lastly, we employ part-of-speech (POS) tagging to distinguish between image-related and language-related tokens. However, relying solely on POS tagging for this differentiation can be problematic. The development of more advanced methods for token distinction could enhance the effectiveness of SumGD and create further synergies with this approach.
@@
In this paper, we contribute to the future development of a safe and reliable AI community by conducting research focused on reducing hallucinations in Large Vision-Language Models Xie et al. (2024); Manevich and Tsarfaty (2024).
@@
However, the models we used may still exhibit biases inherent in the underlying datasets and training processes (Howard et al., 2024; Kim et al., 2024a; Fraser and Kiritchenko, 2024; Lee et al., 2024a; Koh et al., 2024). While our focus was on biases related to language priors, we acknowledge the need to address other potential biases as well.
Acknowledgments
This work was partly supported by the Institute of Information & Communications Technology Planning & Evaluation(IITP)-ITRC(Information Technology Research Center) grant funded by the Korea government(MSIT)(IITP-2025-RS-2024-00437633, 60%), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)] & RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)], and the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2024. K. Jung is with ASRI, Seoul National University, Korea. The Institute of Engineering Research at Seoul National University provided research facilities for this work.
@@ Appendix A Details of Code, Hyperparameters, and GPU Cost
We conduct our experiments based on the OPERA Huang et al. (2024) code base which is publicly available. We use the publicly available code provided by the authors for the VCD and OPERA methods. While M3ID and ICD are implemented from scratch due to the lack of public code. For VCD, OPERA, and ICD, we use the hyperparameters as specified in their respective papers. Since only LLAVA 1.5’s hyperparameters were reported in M3ID, we apply these hyperparameters to both LLAVA 1.5 and InstructBLIP for our experiments. Also, we set repetition penalty as 1. All the experiments are conducted using 1 NVIDIA RTX A6000 GPU.
@@ Appendix B Experimental Settings for Analyzing Language Priors
We generate descriptions using LLAVA 1.5 7B for 5000 images from the MSCOCO 2014 validation dataset Lin et al. (2015) and annotate each token to determine whether it represents an object hallucination. An object hallucination is defined as an object not present in the image. We employ the CHAIR metric pipeline Rohrbach et al. (2019) for evaluation.
Appendix C Ablation Study of Inference Time Cost
Method $RIC$ $C_{s}$ $C_{i}$ $R$
Greedy 1 51.5 13.7 79.1
VCD 2 58.0 16.4 77.8
Beam Search 5 47.5 12.5 79.2
OPERA 5$\uparrow$ 46.0 13.4 78.3
SumGD-S
   +Summarization 2.54 - - -
   +Summarization + POS Tagging 2.98 43.5 11.0 79.2
SumGD-D
   +Summarization 1.87 - - -
   +Summarization + POS Tagging 2.3 42.5 11.6 77.7
Table 5: Comparison of Methods with Relative Inference Costs and CHAIR Metrics. Denote CHAIRS as $C_{S}$, CHAIRI as $C_{I}$, Recall as $R$ and Relative Inference Costs as $RIC$.
We conduct an ablation study to evaluate the impact of the summarization process on inference time. Since the summarization process inherently requires additional token generation, which affects efficiency, we measure inference costs (normalized by relative token generation costs) across baseline methods using the LLAVA 1.5 7B model. Our results show that the distilled summary model (Flan-T5-base) in SumGD-D requires only half the time per token compared to the LLAVA 1.5 7B model. This finding is incorporated into our inference cost calculations. Additionally, to ensure accurate part-of-speech (POS) tagging, we generate one extra word after each current token and then measure the current token’s POS tag. We perform this ablation study on 200 images. The comparison of inference time costs is presented in Table 5.
Appendix D Experiments on LLaVA 1.6
Method CHAIRS CHAIRI
Greedy 36.5 10.2
Nucleus 37.0 9.7
Beam-Search (beams=5) 34.5 10.5
SumGD-S 30.5 6.4
Table 6: CHAIR results on LLAVA 1.6 7B (max new tokens is 512).
We conduct an evaluation of our SumGD method on the latest model, LLaVA 1.6 7B model Liu et al. (2024d), to assess its effectiveness. Specifically, we perform the CHAIR evaluation (lower is better) on 200 images. Table 6 shows that SumGD-S effectively reduces object hallucination, demonstrating the applicability of our methodology even with the latest models.
Appendix E Distilled Flan-T5-base model
We employ LLAVA 1.5 7B to perform Summary-Guided Decoding with Self-Summarization while generating descriptions for 5,000 images from the MSCOCO dataset. During this process, LLAVA 1.5 iteratively summarizes each previous sentence, and we store each previous sentence along with its corresponding summarized sentence as a pair. This paired dataset is subsequently used to fine-tune the Flan-T5-base model with the prompt “What is a summary of this text?” for training purposes.
Appendix F Summarize Prompt for Summary-Guided Decoding
In SumGD-S, we use summary prompt as:
USER: Summarize the following
@@ Appendix F Summarize Prompt for Summary-Guided Decoding
In SumGD-D, we use summary prompt as:
[…]
this text?
Appendix G GPT-4o Prompt for text quality evaluation
@@
Appendix H CHAIR metric on various token length
@@
Appendix I Case Study
This is the case study of Summary-Guided Decoding (SumGD), Visual Contrastive Decoding (VCD), and Multi-Modal Mutual Information Decoding (M3ID) in generating up to 256 tokens in detailed captioning task. Case study shows that SumGD generated image-related words, while VCD showed a tendency to hallucinate by relying on the word ‘tie’ during decoding. Additionally, M3ID exhibited issues in language modeling.
Figure 9: SumGD case study.